# 1S-DAug: One-Shot Data Augmentation for Robust Few-Shot Generalization

### v1

@@ B.3 Equivalence of Logit Averaging and Feature Averaging
$\bar{z}(q):=\sum_{k}\alpha_{k}\,z^{(k)}(q),\qquad g_{\theta}^{\mathrm{avg}}(q,p):=-\bigl\|\bar{z}(q)-p\bigr\|_{2}^{2}.$
@@
$\tilde{g}_{\theta}(q,p)$ $=-\sum_{k}\alpha_{k}\bigl\|z^{(k)}(q)-p\bigr\|_{2}^{2}$
$=-\sum_{k}\alpha_{k}\Bigl(\|z^{(k)}(q)\|_{2}^{2}-2\langle z^{(k)}(q),p\rangle+\|p\|_{2}^{2}\Bigr)$
$=-\sum_{k}\alpha_{k}\|z^{(k)}(q)\|_{2}^{2}+2\Bigl\langle\sum_{k}\alpha_{k}z^{(k)}(q),p\Bigr\rangle-\|p\|_{2}^{2}$
$=-\sum_{k}\alpha_{k}\|z^{(k)}(q)\|_{2}^{2}-\|\bar{z}(q)\|_{2}^{2}+\|\bar{z}(q)\|_{2}^{2}+2\langle\bar{z}(q),p\rangle-\|p\|_{2}^{2}$
$=-\bigl\|\bar{z}(q)-p\bigr\|_{2}^{2}-\sum_{k}\alpha_{k}\|z^{(k)}(q)\|_{2}^{2}+\|\bar{z}(q)\|_{2}^{2}.$
The last two terms depend on $q$ and the set of views $\{z^{(k)}(q)\}$ but not on $p$. Hence, for fixed $q$, comparing classes by $\tilde{g}_{\theta}(q,p_{c})$ is equivalent to comparing them by
$g_{\theta}^{\mathrm{avg}}(q,p_{c})=-\bigl\|\bar{z}(q)-p_{c}\bigr\|_{2}^{2}.$
In other words, query-side logit averaging with squared Euclidean scores induces exactly the same class ranking as nearest-prototype classification in the feature space of the averaged query embedding $\bar{z}(q)$.
This justifies working, for analysis, with the aggregated difference feature
$\bar{\Omega}_{\theta}(x):=\bar{z}(q)-p$
and its radius, even though the implementation performs logit averaging rather than explicit feature averaging on the query side.
B.4 Stability of the Empirical Margin Risk
Let $S=\{(x_{i},y_{i})\}_{i=1}^{m}$ be the original sample and let $\widehat{S}$ denote the same pairs $(x_{i},y_{i})$ but with the empirical margin risk evaluated using the aggregated difference features $\bar{\Omega}_{\theta}(x_{i})$. Write
$\Omega_{i}:=\Omega_{\theta}(x_{i}),\qquad\widehat{\Omega}_{i}:=\bar{\Omega}_{\theta}(x_{i}),$
and recall that $g_{\theta}(x)=-\|\Omega_{\theta}(x)\|_{2}^{2}$. Abusing notation slightly, we write $g_{\theta}(\Omega)$ for the score evaluated at a difference feature $\Omega$, i.e. $g_{\theta}(\Omega):=-\|\Omega\|_{2}^{2}$. Assume that $g_{\theta}(\cdot)$ is $L_{g}$-Lipschitz with respect to its feature argument on the radius-$r_{0}$ ball in $\mathbb{R}^{d}$.
Lemma B.3 (Empirical margin stability).
For any $\rho>0$,
$\bigl|\widehat{R}_{\widehat{S},\rho}(\theta)-\widehat{R}_{S,\rho}(\theta)\bigr|\;\leq\;\frac{L_{g}}{\rho m}\sum_{i=1}^{m}\|\widehat{\Omega}_{i}-\Omega_{i}\|_{2}.$ (18)
In particular,
$\bigl|\widehat{R}_{\widehat{S},\rho}(\theta)-\widehat{R}_{S,\rho}(\theta)\bigr|\;\leq\;\frac{L_{g}}{\rho}\cdot\Bigl(\frac{1}{m}\sum_{i=1}^{m}\|\widehat{\Omega}_{i}-\Omega_{i}\|_{2}\Bigr).$ (19)
@@ B.4 Stability of the Empirical Margin Risk
For each $i$,
$\bigl|\tau_{\rho}\bigl(y_{i}g_{\theta}(\widehat{\Omega}_{i})\bigr)-\tau_{\rho}\bigl(y_{i}g_{\theta}(\Omega_{i})\bigr)\bigr|\leq\frac{1}{\rho}\,\bigl|y_{i}g_{\theta}(\widehat{\Omega}_{i})-y_{i}g_{\theta}(\Omega_{i})\bigr|=\frac{1}{\rho}\,\bigl|g_{\theta}(\widehat{\Omega}_{i})-g_{\theta}(\Omega_{i})\bigr|.$
Since $g_{\theta}$ is $L_{g}$-Lipschitz in its feature argument,
$\bigl|g_{\theta}(\widehat{\Omega}_{i})-g_{\theta}(\Omega_{i})\bigr|\leq L_{g}\,\|\widehat{\Omega}_{i}-\Omega_{i}\|_{2}.$
Combining the two displays gives
$\bigl|\tau_{\rho}\bigl(y_{i}g_{\theta}(\widehat{\Omega}_{i})\bigr)-\tau_{\rho}\bigl(y_{i}g_{\theta}(\Omega_{i})\bigr)\bigr|\leq\frac{L_{g}}{\rho}\|\widehat{\Omega}_{i}-\Omega_{i}\|_{2}.$
Averaging over $i$ and using the triangle inequality yields (18). ∎
For the Euclidean score class considered here we can take $L_{g}\leq 2r_{0}$, since $z\mapsto-\|z\|_{2}^{2}$ has gradient of norm $2\|z\|_{2}$ and is therefore $2r_{0}$-Lipschitz on the radius-$r_{0}$ ball.
B.5 Probabilistic Radius Reduction with Multiple Augmentations
For each $i=1,\dots,m$ and $k=0,\dots,M$, let $\{\Omega_{i}^{(k)}\in\mathbb{R}^{d}\}$ be i.i.d. feature vectors with radii $R_{i}^{(k)}:=\|\Omega_{i}^{(k)}\|_{2}$. Define
$R_{\mathrm{orig}}^{\max}:=\max_{1\leq i\leq m}R_{i}^{(0)},\qquad R_{\mathrm{aug}}^{\max}:=\max_{1\leq i\leq m}\max_{1\leq k\leq M}R_{i}^{(k)}.$ (20)
Define the aggregated feature and corresponding radii
$\widehat{\Omega}_{i}:=\frac{1}{M+1}\sum_{k=0}^{M}\Omega_{i}^{(k)},\quad\widehat{R}_{i}:=\|\widehat{\Omega}_{i}\|_{2},\quad\widehat{R}^{\max}:=\max_{1\leq i\leq m}\widehat{R}_{i}.$ (21)
Proposition B.4 (Radius reduction with $M$ augmentations).
Assume $\{\Omega_{i}^{(k)}\}_{i,k}$ are i.i.d. from a continuous distribution on $\mathbb{R}^{d}$. Then
$\mathbb{P}\bigl(\widehat{R}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)\;\geq\;\mathbb{P}\bigl(R_{\mathrm{aug}}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)=\frac{1}{M+1}.$ (22)
Proof.
Consider all $m(M+1)$ radii $\{R_{i}^{(k)}\}_{i,k}$, which are i.i.d. on $\mathbb{R}_{+}$. Let $R_{\mathrm{all}}^{\max}:=\max_{i,k}R_{i}^{(k)}$. By continuity, this maximum is almost surely unique and, by symmetry, each of the $m(M+1)$ radii is equally likely to be the maximum. The event $\{R_{\mathrm{aug}}^{\max}<R_{\mathrm{orig}}^{\max}\}$ occurs exactly when the unique maximum lies among the $m$ original radii $\{R_{i}^{(0)}\}$, hence
$\mathbb{P}\bigl(R_{\mathrm{aug}}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)=\frac{m}{m(M+1)}=\frac{1}{M+1}.$
Now fix an outcome in this event. Let $I$ be the (unique) index such that $R_{I}^{(0)}=R_{\mathrm{orig}}^{\max}$. Then for every $i$ and $k\geq 1$, $R_{i}^{(k)}<R_{\mathrm{orig}}^{\max}$, and for all $j\neq I$, $R_{j}^{(0)}<R_{\mathrm{orig}}^{\max}$. For $i=I$, the average of the $M+1$ vectors $\{\Omega_{I}^{(k)}\}$ has strictly smaller norm than the largest of them: not all $\Omega_{I}^{(k)}$ are equal (by continuity), and the Euclidean norm is strictly convex, so
$\widehat{R}_{I}=\biggl\|\frac{1}{M+1}\sum_{k=0}^{M}\Omega_{I}^{(k)}\biggr\|_{2}<R_{\mathrm{orig}}^{\max}.$
For any $j\neq I$, each $\|\Omega_{j}^{(k)}\|_{2}$ is strictly less than $R_{\mathrm{orig}}^{\max}$, hence by the triangle inequality
$\widehat{R}_{j}=\biggl\|\frac{1}{M+1}\sum_{k=0}^{M}\Omega_{j}^{(k)}\biggr\|_{2}\leq\frac{1}{M+1}\sum_{k=0}^{M}\|\Omega_{j}^{(k)}\|_{2}<R_{\mathrm{orig}}^{\max}.$
Thus $\widehat{R}^{\max}<R_{\mathrm{orig}}^{\max}$ on this event, so
$\mathbb{P}\bigl(\widehat{R}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)\geq\mathbb{P}\bigl(R_{\mathrm{aug}}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)=\frac{1}{M+1}.$
B.6 Training-time versus Test-time Augmentation

### v3

@@ B.3 Margin-based Bound and Rademacher Complexity
$\mathcal{H}_{\mathrm{enc}}:=\bigl\{e_{\theta}:q\mapsto\Phi_{\theta}(q)\;:\;\theta\in\Theta\bigr\}\subseteq\mathbb{R}^{d}.$
Let
$\varphi(u):=\beta-\|u\|_{2}^{2},\qquad u\in\mathbb{R}^{d},$
where $\beta\in\mathbb{R}$ is the fixed threshold from the score definition. Then $g_{\theta}(x)=\varphi(\Omega_{\theta}(x))$ for every $\theta$, so $\mathcal{G}=\varphi\circ\mathcal{F}$.
We first control the Lipschitz constant of the outer map $\varphi$ on the relevant domain. Fix $u,v\in\mathbb{R}^{d}$ with $\|u\|_{2}\leq r$ and $\|v\|_{2}\leq r$. Then
$\begin{split}|\varphi(u)-\varphi(v)|&=\bigl|\,\|v\|_{2}^{2}-\|u\|_{2}^{2}\,\bigr|\\ &=\bigl|\langle u+v,\,u-v\rangle\bigr|\\ &\leq\|u+v\|_{2}\,\|u-v\|_{2}\\ &\leq(\|u\|_{2}+\|v\|_{2})\,\|u-v\|_{2}\\ &\leq 2r\,\|u-v\|_{2}.\end{split}$ (23)
Hence $\varphi$ is $2r$-Lipschitz on the radius-$r$ ball. Since $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$ by assumption, (23) applies to every value attained by every $f\in\mathcal{F}$ on the sample.
Next, define the empirical vector Rademacher complexity of a class $\mathcal{V}\subseteq(\mathbb{R}^{d})^{\mathcal{X}}$ on a sample $\mathcal{B}=\{b_{i}\}_{i=1}^{m}$ by
$\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{B}}(\mathcal{V}):=\frac{1}{m}\,\mathbb{E}_{\xi}\left[\sup_{f\in\mathcal{V}}\sum_{i=1}^{m}\sum_{\mu=1}^{d}\xi_{i\mu}\,[f(b_{i})]_{\mu}\right],$
where $\{\xi_{i\mu}\}$ are i.i.d. Rademacher variables. By the vector contraction inequality, there exists a universal constant $C_{0}>0$ such that
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})=\widehat{\mathfrak{R}}_{\mathcal{T}}(\varphi\circ\mathcal{F})\leq C_{0}\,\operatorname{Lip}(\varphi)\,\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{T}}(\mathcal{F}).$ (24)
Using $\operatorname{Lip}(\varphi)\leq 2r$ on the relevant domain, we obtain
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})\leq 2C_{0}r\,\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{T}}(\mathcal{F}).$ (25)
We now remove the prototype term. Since $\Omega_{\theta}(x_{i})=\Phi_{\theta}(q_{i})-p_{i}$, we have
$\begin{split}\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{T}}(\mathcal{F})&=\frac{1}{m}\,\mathbb{E}_{\xi}\left[\sup_{\theta\in\Theta}\sum_{i=1}^{m}\sum_{\mu=1}^{d}\xi_{i\mu}\bigl([\Phi_{\theta}(q_{i})]_{\mu}-[p_{i}]_{\mu}\bigr)\right]\\ &=\frac{1}{m}\,\mathbb{E}_{\xi}\left[\sup_{\theta\in\Theta}\sum_{i=1}^{m}\sum_{\mu=1}^{d}\xi_{i\mu}[\Phi_{\theta}(q_{i})]_{\mu}\right]-\frac{1}{m}\,\mathbb{E}_{\xi}\left[\sum_{i=1}^{m}\sum_{\mu=1}^{d}\xi_{i\mu}[p_{i}]_{\mu}\right]\\ &=\frac{1}{m}\,\mathbb{E}_{\xi}\left[\sup_{\theta\in\Theta}\sum_{i=1}^{m}\sum_{\mu=1}^{d}\xi_{i\mu}[\Phi_{\theta}(q_{i})]_{\mu}\right]\\ &=\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{Q}_{\mathcal{T}}}(\mathcal{H}_{\mathrm{enc}}),\end{split}$ (26)
where $\mathcal{Q}_{\mathcal{T}}:=\{q_{i}\}_{i=1}^{m}$ and the second term vanishes because $\mathbb{E}[\xi_{i\mu}]=0$.
It remains to bound $\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{Q}_{\mathcal{T}}}(\mathcal{H}_{\mathrm{enc}})$ by a covering-number argument specialized to the present architecture. Write the encoder as
$\Phi_{\theta}=\Phi_{W_{1},\dots,W_{L}},$
where layer $\ell$ has width $n_{\ell}$ (so $n_{L}=d$),
$a_{0}(q):=q,\qquad a_{\ell}(q):=\psi_{\ell}\bigl(W_{\ell}a_{\ell-1}(q)\bigr),\qquad\Phi_{\theta}(q)=a_{L}(q),$
and each nonlinearity $\psi_{\ell}$ is $1$-Lipschitz. Since the input domain and preprocessing are fixed throughout the paper, there exists a finite radius $B_{X}:=\sup_{q\in\mathcal{X}}\|q\|_{2}<\infty$.
We first record the uniform layerwise output bound
$\|a_{\ell}(q)\|_{2}\leq\Bigl(\prod_{\lambda=1}^{\ell}s_{\lambda}\Bigr)\,\|q\|_{2}\leq B_{X}\,\prod_{\lambda=1}^{\ell}s_{\lambda}\qquad\text{for all }q\in\mathcal{X}\text{ and }\ell=1,\dots,L.$ (27)
Indeed, this follows inductively from $\|a_{\ell}(q)\|_{2}\leq\|W_{\ell}\|_{2}\,\|a_{\ell-1}(q)\|_{2}\leq s_{\ell}\,\|a_{\ell-1}(q)\|_{2}$. In particular, for the encoder output one has
$\|\Phi_{\theta}(q)\|_{2}=\|a_{L}(q)\|_{2}\leq B_{X}\prod_{\lambda=1}^{L}s_{\lambda}\leq B_{X}L_{\mathrm{enc}}.$ (28)
Next we control how the encoder changes under perturbations of the layer weights. Let $\theta=(W_{1},\dots,W_{L})$ and $\theta^{\prime}=(W_{1}^{\prime},\dots,W_{L}^{\prime})$, and write $a_{\ell}$ and $a_{\ell}^{\prime}$ for the corresponding hidden states. For each $q\in\mathcal{X}$,
$\begin{split}\|a_{\ell}(q)-a_{\ell}^{\prime}(q)\|_{2}&\leq\bigl\|W_{\ell}a_{\ell-1}(q)-W_{\ell}^{\prime}a_{\ell-1}^{\prime}(q)\bigr\|_{2}\\ &\leq\bigl\|W_{\ell}\bigl(a_{\ell-1}(q)-a_{\ell-1}^{\prime}(q)\bigr)\bigr\|_{2}+\bigl\|(W_{\ell}-W_{\ell}^{\prime})a_{\ell-1}^{\prime}(q)\bigr\|_{2}\\ &\leq s_{\ell}\,\|a_{\ell-1}(q)-a_{\ell-1}^{\prime}(q)\|_{2}+\|W_{\ell}-W_{\ell}^{\prime}\|_{2}\,\|a_{\ell-1}^{\prime}(q)\|_{2}\\ &\leq s_{\ell}\,\|a_{\ell-1}(q)-a_{\ell-1}^{\prime}(q)\|_{2}+B_{X}\Bigl(\prod_{\lambda<\ell}s_{\lambda}\Bigr)\|W_{\ell}-W_{\ell}^{\prime}\|_{2},\end{split}$ (29)
where the last line uses (27) for $a_{\ell-1}^{\prime}$. Since $a_{0}(q)=a_{0}^{\prime}(q)=q$, repeatedly unrolling (29) gives
$\begin{split}\|\Phi_{\theta}(q)-\Phi_{\theta^{\prime}}(q)\|_{2}&\leq B_{X}\sum_{\ell=1}^{L}\Bigl(\prod_{\lambda<\ell}s_{\lambda}\Bigr)\Bigl(\prod_{\lambda>\ell}s_{\lambda}\Bigr)\|W_{\ell}-W_{\ell}^{\prime}\|_{2}\\ &\leq B_{X}L_{\mathrm{enc}}\sum_{\ell=1}^{L}\frac{\|W_{\ell}-W_{\ell}^{\prime}\|_{2}}{s_{\ell}}.\end{split}$ (30)
Therefore the same bound holds for the empirical metric
$\|e_{\theta}-e_{\theta^{\prime}}\|_{\mathcal{Q}_{\mathcal{T}},2}:=\Bigl(\frac{1}{m}\sum_{i=1}^{m}\|\Phi_{\theta}(q_{i})-\Phi_{\theta^{\prime}}(q_{i})\|_{2}^{2}\Bigr)^{1/2},$
since pointwise control implies control of the sample average.
For each layer define the spectral-norm ball
$\mathcal{W}_{\ell}:=\bigl\{W\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}:\ \|W\|_{2}\leq s_{\ell}\bigr\}.$
If $s_{\ell}=0$ for some $\ell$, then $\mathcal{W}_{\ell}=\{0\}$ has covering number $1$, and that layer may be omitted from the covering-number calculation. Thus, without loss of generality, we may assume $s_{\ell}>0$ for all $\ell$ below. Fix $\gamma>0$, and choose layerwise accuracies
$\eta_{\ell}:=\frac{\gamma s_{\ell}}{B_{X}LL_{\mathrm{enc}}},\qquad\ell=1,\dots,L.$ (31)
If $\|W_{\ell}-W_{\ell}^{\prime}\|_{2}\leq\eta_{\ell}$ for every $\ell$, then by (30) we obtain
$\|e_{\theta}-e_{\theta^{\prime}}\|_{\mathcal{Q}_{\mathcal{T}},2}\leq B_{X}L_{\mathrm{enc}}\sum_{\ell=1}^{L}\frac{\eta_{\ell}}{s_{\ell}}=\gamma.$ (32)
Hence an $\eta_{\ell}$-cover of each $\mathcal{W}_{\ell}$ induces a $\gamma$-cover of $\mathcal{H}_{\mathrm{enc}}$ on the sample $\mathcal{Q}_{\mathcal{T}}$.
We now bound the size of those layerwise covers. By the volumetric estimate for a radius-$s_{\ell}$ ball in the $n_{\ell}n_{\ell-1}$-dimensional matrix space,
$N\bigl(\eta_{\ell},\mathcal{W}_{\ell},\|\cdot\|_{2}\bigr)\leq\Bigl(1+\frac{2s_{\ell}}{\eta_{\ell}}\Bigr)^{n_{\ell}n_{\ell-1}}.$ (33)
Let
$N_{\mathrm{par}}:=\sum_{\ell=1}^{L}n_{\ell}n_{\ell-1}$
be the total number of scalar weight parameters. We now make the passage from layerwise covers to the encoder-class cover explicit. For each $\ell$, let $\mathcal{C}_{\ell}$ be an $\eta_{\ell}$-cover of $\mathcal{W}_{\ell}$ in spectral norm, with cardinality
$|\mathcal{C}_{\ell}|=N\bigl(\eta_{\ell},\mathcal{W}_{\ell},\|\cdot\|_{2}\bigr).$
Then every parameter tuple $(W_{1},\dots,W_{L})\in\prod_{\ell=1}^{L}\mathcal{W}_{\ell}$ can be approximated by a tuple $(\widetilde{W}_{1},\dots,\widetilde{W}_{L})\in\prod_{\ell=1}^{L}\mathcal{C}_{\ell}$ such that $\|W_{\ell}-\widetilde{W}_{\ell}\|_{2}\leq\eta_{\ell}$ for all $\ell$. By the previous paragraph, the corresponding encoders satisfy
$\|e_{\theta}-e_{\widetilde{\theta}}\|_{\mathcal{Q}_{\mathcal{T}},2}\leq\gamma.$
Hence
$N\bigl(\gamma,\mathcal{H}_{\mathrm{enc}}|_{\mathcal{Q}_{\mathcal{T}}},\|\cdot\|_{\mathcal{Q}_{\mathcal{T}},2}\bigr)\leq\prod_{\ell=1}^{L}N\bigl(\eta_{\ell},\mathcal{W}_{\ell},\|\cdot\|_{2}\bigr).$
Applying (33) layerwise gives
$\begin{split}N\bigl(\gamma,\mathcal{H}_{\mathrm{enc}}|_{\mathcal{Q}_{\mathcal{T}}},\|\cdot\|_{\mathcal{Q}_{\mathcal{T}},2}\bigr)&\leq\prod_{\ell=1}^{L}\Bigl(1+\frac{2s_{\ell}}{\eta_{\ell}}\Bigr)^{n_{\ell}n_{\ell-1}}\\ &=\prod_{\ell=1}^{L}\Bigl(1+\frac{2B_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)^{n_{\ell}n_{\ell-1}}\\ &=\Bigl(1+\frac{2B_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)^{\sum_{\ell=1}^{L}n_{\ell}n_{\ell-1}}\\ &=\Bigl(1+\frac{2B_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)^{N_{\mathrm{par}}},\end{split}$ (34)
where in the second line we used $\eta_{\ell}=\gamma s_{\ell}/(B_{X}LL_{\mathrm{enc}})$ from (31), so the ratio $2s_{\ell}/\eta_{\ell}$ is independent of $\ell$. Taking logarithms yields
$\begin{split}\log N\bigl(\gamma,\mathcal{H}_{\mathrm{enc}}|_{\mathcal{Q}_{\mathcal{T}}},\|\cdot\|_{\mathcal{Q}_{\mathcal{T}},2}\bigr)&\leq N_{\mathrm{par}}\,\log\Bigl(1+\frac{2B_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)\\ &\leq N_{\mathrm{par}}\,\log\Bigl(\frac{CB_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)\end{split}$ (35)
for a universal constant $C>0$. In the range used later in Dudley’s integral, namely $0<\gamma\leq 2B_{X}L_{\mathrm{enc}}$ by (36), we have $2B_{X}LL_{\mathrm{enc}}/\gamma\geq L\geq 1$, so one may take, for example, $C=4$ since $1+t\leq 2t$ for $t\geq 1$.
We next bound the diameter of the sample-restricted encoder class. By (28),
$\|\Phi_{\theta}(q_{i})\|_{2}\leq B_{X}L_{\mathrm{enc}}\qquad\text{for every }i=1,\dots,m,$
so
$\operatorname{diam}\bigl(\mathcal{H}_{\mathrm{enc}}|_{\mathcal{Q}_{\mathcal{T}}},\|\cdot\|_{\mathcal{Q}_{\mathcal{T}},2}\bigr)\leq 2B_{X}L_{\mathrm{enc}}.$ (36)
Applying Dudley’s entropy integral in the vector-valued form gives a universal constant $C_{1}>0$ such that
$\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{Q}_{\mathcal{T}}}(\mathcal{H}_{\mathrm{enc}})\leq\frac{C_{1}}{\sqrt{m}}\int_{0}^{2B_{X}L_{\mathrm{enc}}}\sqrt{\log N\bigl(\gamma,\mathcal{H}_{\mathrm{enc}}|_{\mathcal{Q}_{\mathcal{T}}},\|\cdot\|_{\mathcal{Q}_{\mathcal{T}},2}\bigr)}\,d\gamma.$ (37)
Substituting (35) into (37) yields
$\begin{split}\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{Q}_{\mathcal{T}}}(\mathcal{H}_{\mathrm{enc}})&\leq\frac{C_{1}\sqrt{N_{\mathrm{par}}}}{\sqrt{m}}\int_{0}^{2B_{X}L_{\mathrm{enc}}}\sqrt{\log\Bigl(\frac{CB_{X}LL_{\mathrm{enc}}}{\gamma}\Bigr)}\,d\gamma\\ &\leq\frac{C_{2}B_{X}L_{\mathrm{enc}}\sqrt{N_{\mathrm{par}}}}{\sqrt{m}}\end{split}$ (38)
for another universal constant $C_{2}>0$, where the last step follows from the change of variables $u=\log\bigl(CB_{X}LL_{\mathrm{enc}}/\gamma\bigr)$, which shows that the integral is bounded by a constant multiple of $B_{X}L_{\mathrm{enc}}$.
Since the input dimension and widths $(n_{0},\dots,n_{L})$ are fixed throughout the paper, the factor $\sqrt{N_{\mathrm{par}}}$ may be absorbed into a constant depending only on the architecture. Therefore there exists $C_{\mathrm{arch}}>0$, depending only on the architecture, such that
$\widehat{\mathfrak{R}}^{\mathrm{vec}}_{\mathcal{Q}_{\mathcal{T}}}(\mathcal{H}_{\mathrm{enc}})\leq C_{\mathrm{arch}}\,B_{X}\,\frac{L_{\mathrm{enc}}}{\sqrt{m}}.$ (39)
Combining (25), (26), and (39), we obtain
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})\leq 2C_{0}C_{\mathrm{arch}}\,B_{X}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}}.$
Therefore the claim follows by setting
$C_{\mathrm{enc}}:=2C_{0}C_{\mathrm{arch}}.$
The final statement follows by substituting the corresponding radius bound: $r=r_{0}$ before augmentation and $r=\hat{r}$ after augmentation. ∎
B.4 Probabilistic Radius Reduction under Augmentation
Proposition B.3 (Radius reduction with one augmentation).
Assume $\{\Omega_{i}^{(0)},\Omega_{i}^{(1)}\}_{i=1}^{m}$ are i.i.d. from a continuous distribution on $\mathbb{R}^{d}$. Let
$\bar{\Omega}_{i}:=\frac{1}{2}\bigl(\Omega_{i}^{(0)}+\Omega_{i}^{(1)}\bigr),\;\hat{r}^{\max}:=\max_{i}\|\bar{\Omega}_{i}\|_{2},\;r^{\max}_{\mathrm{aug}}:=\max_{i}\|{\Omega}_{i}^{(1)}\|_{2},\;r_{\mathrm{orig}}^{\max}:=\max_{i}\|\Omega_{i}^{(0)}\|_{2}.$ (40)
@@ B.4 Probabilistic Radius Reduction under Augmentation
$\mathbb{P}\bigl(\hat{r}^{\max}<r_{\mathrm{orig}}^{\max}\bigr)>0.5.$ (41)
@@
Since $\{\Omega_{i}^{(0)}\}_{i=1}^{m}$ and $\{\Omega_{i}^{(1)}\}_{i=1}^{m}$ are i.i.d. and the distribution is continuous, the random variables $r^{\max}_{\mathrm{orig}}$ and $r^{\max}_{\mathrm{aug}}$ are i.i.d. and, by symmetry,
$\mathbb{P}(r^{\max}_{\mathrm{aug}}<r^{\max}_{\mathrm{orig}})=\mathbb{P}(r^{\max}_{\mathrm{orig}}<r^{\max}_{\mathrm{aug}})=\tfrac{1}{2}.$
Case 1: On the event $r^{\max}_{\mathrm{aug}}<r^{\max}_{\mathrm{orig}}$, we have $\|\Omega_{i}^{(1)}\|_{2}<r^{\max}_{\mathrm{orig}}$ for every $i$, while by definition $\|\Omega_{i}^{(0)}\|_{2}\leq r^{\max}_{\mathrm{orig}}$ for every $i$. Therefore, for every $i$,
$\|\bar{\Omega}_{i}\|_{2}=\left\|\frac{\Omega_{i}^{(0)}+\Omega_{i}^{(1)}}{2}\right\|_{2}\leq\frac{\|\Omega_{i}^{(0)}\|_{2}+\|\Omega_{i}^{(1)}\|_{2}}{2}<r^{\max}_{\mathrm{orig}},$
where the inequality uses the triangle inequality.
Case 2: When the new augmentation set maximum feature radius is larger, there are possible events that either increase or decrease the aggregated feature radius under this case; the augmentation set maximum feature radius may end up pairing a small-magnitude original feature radius, leading to decreased maximum feature radius after the averaging. Since that the maximum feature radius decreases for sure in Case 1, and does not increase for sure in Case 2, the probability of feature radius increase after the averaging is less than 0.5. ∎
Remark B.4.
When there is more than one augmentation set, we can treat the aggregated set as the augmentation set altogether, and the probability of maximum feature radius increase becomes even lower due to the expected variance reduction (i.e., extremely large feature radius) within the augmentation set.
Remark B.5.
The same feature radius reduction is for the feature difference $\Omega$, but is also applicable to the single query feature (i.e., $B_{X}$).
B.5 Training-time versus Test-time Augmentation